two identicle baals each having a density rho r suspended frm a commn pt by two insulated copper stringsof eql length . both the balls  have eq mass nd charge . in equllibrium both of them frm an angle theta frm vert nw they r immeresed in liquid of densit sigma (as the theta do not changes in liquid)find the di electric cons tant of the liquid

Grade:12

2 Answers

aakanksha vyas
12 years ago

sigma/ sigma-rho=dielectric constant

Chetan Mandayam Nayakar
12 years ago

let 't' be tension in string. tcosθ=mg, tsinθ=(q2/4πε0r2) r is distance between two balls.

t(1-(σ/ρ))=(q2/4πε0kr2), k is the dielectric constant